### v1

# Bilayer graphene quantum dots as a quantum simulator of Haldane topological quantum matter

Abstract
We demonstrate here that a chain of Bilayer Graphene Quantum Dots (BLGQD) can realize topological quantum matter by effectively simulating a spin-1 chain that hosts the Haldane phase within a specific range of parameters. We describe a chain of BLGQD with two electrons each using an atomistic tight-binding model combined with the exact diagonalization technique to solve the interacting few-electron problem. Coulomb interactions and valley mixing effects are treated within the same microscopic framework, allowing us to systematically investigate spin and valley polarization transitions as functions of interaction strength and external tuning parameters. We calculate the low energy states for single and double QDs as a function of the number of electrons, identifying regimes of highly correlated multi-electron states. We confirm the presence of a spin-one ground state for two electrons. Then, we explore two coupled QDs with 4 electrons and extend the analysis to QD arrays. Using a mapping of the BLGQD chain to an effective bilinear-biquadratic (BLBQ) spin model, we demonstrate that BLGQD arrays can work as a quantum simulator for one-dimensional spin chains with emergent many-body topological phases.
Advancing quantum information technologies and uncovering novel topological phases of matter are central goals in modern condensed matter physics and quantum engineering [1, 2, 3, 4]. A primary focus is the development of robust qubits with extended coherence times, with various platforms, including uperconducting [5, 6, 7, 8], trapped ion [9, 10, 11], photonic [12, 13, 14, 15], and semiconductor spin qubits [16, 17, 3, 18, 19, 20, 21, 22, 23, 24, 25, 26]. Spin qubits and circuits, chains, stand out as prototypes for investigating topologically strongly correlated quantum matter, particularly due to their capacity to host Haldane spin-1/2 quasiparticles at their edges [27, 28, 29, 30].
The artificial synthetic spin-1 chains have been intensely studied, but their experimental realization is limited. Spin-1 chains can be engineered using various quantum dot architectures, such as gated triple quantum dots [31], linear arrays of semiconductor quantum dots embedded in nanowires [32, 30, 33, 34], chains formed by triangular graphene quantum dots [35, 36, 37, 38], or tunable hybrid platform of superconducting islands and quantum dots [39]. For instance, chains of InAsP quantum dots embedded in an InP nanowire have been proposed and predicted to host such synthetic spin-1 objects [33, 40], realizing macroscopic quantum states in semiconductors. In these systems, four electrons per quantum dot can form a synthetic spin-1 state, and their low-energy behaviour can be effectively described by a Hubbard-Kanamori Hamiltonian derived from atomistic microscopic calculations [40]. This framework has successfully demonstrated that these arrays can emulate antiferromagnetic spin-1 chains, providing a foundation for engineering synthetic topologically nontrivial quantum matter. Gated bilayer graphene (BLG) quantum dots (QDs) offer a particularly attractive platform for this purpose. Single and double BGQDs have been experimentally realized, with evidence of spin-1 ground states in few-electron configurations [41, 42]. BLG is unique among two-dimensional materials in that a perpendicular electric field can open a tunable band gap, enabling electrical control over confinement and access to both spin and valley degrees of freedom [41, 43, 42]. Recent theoretical and experimental studies of BLGQDs show that two confined electrons can robustly form a triplet ground state across a wide range of interaction strengths, often accompanied by spontaneous spin–valley polarization [42, 41]. This robust triplet formation makes BLGQDs promising building blocks for electrically tunable synthetic spin-1 chains.
Here, we investigate single and coupled BLGQDs using an atomistic tight-binding model combined with exact diagonalization to treat Coulomb interactions and valley mixing on equal footing. For two coupled QDs, we show that the low-energy sector maps naturally onto a bilinear–biquadratic spin-1 model, establishing a direct link between microscopic BLGQD states and effective spin-1 chains. This mapping highlights BLGQDs as a practical and tunable platform for exploring one-dimensional quantum magnetism, topological phases, and electrically controlled quantum information architectures.
We model Bernal-stacked bilayer graphene (BLG), Fig.1(a), whose bottom (top) layer contains sublattices $A_{1}$ and $B_{1}$ ($A_{2}$ and $B_{2}$). The electronic structure is described within a tight-binding model including the dominant in-plane and interlayer hopping parameters ($\gamma_{0}$, $\gamma_{1}$) and a perpendicular electric field that opens a tunable band gap [Fig.1(b)]. Gate-defined quantum dots (QDs) are implemented via smooth Gaussian confinement potentials [Fig. 1(c)], enabling either single- or double-dot configurations. Confined states are obtained by projecting onto bulk Bloch states near the $\pm K$ valleys and diagonalizing the Hamiltonian including the external potential. Calculations employ a rhomboidal supercell of $901\times 901$ unit cells ($\sim 3\times 10^{6}$ atoms). Full details of the tight-binding model, confinement treatment in momentum space, and numerical implementation are provided in the Supplementary Material.
Figure 1: (a) Top view of bilayer graphene, showing the four atoms within the unit cell: $A_{1}$, $B_{1}$, $A_{2}$, $B_{2}$, distinguished by colour. (b) Side view (zoomed in), highlighting the two primary hopping parameters: $\gamma_{0}$ for intralayer coupling and $\gamma_{1}$ for interlayer coupling. The applied perpendicular electric field is shown schematically; note that the diagram depicts the resulting potential energy profile rather than the electrostatic potential. (c) Schematic illustration of the confining potential for a double quantum dot. The parameters $V_{0}$, $R_{\rm{QD}}$, and $d$ denote the potential depth, dot radius, and center-to-center distance between the dots, respectively.
Many-body effects are incorporated by filling the lowest single-particle levels with $N$ electrons and adding the Coulomb interaction,
[…]
where $E_{s}$ are the single-particle energies, $\langle pq|V|rs\rangle$ are the Coulomb matrix elements, and $\lambda\in[0,1]$ tunes the interaction strength. The spin structure is included explicitly, and all matrix elements are evaluated using the numerically obtained single-particle orbitals. The resulting many-body Hamiltonian in the space of multi-electron configurations is solved via exact diagonalization to obtain the low-energy spectrum [42, 44].
For completeness, we first characterize the single quantum dot (QD) spectrum as a reference for the analysis of the double QD and BLG QD chain. For a representative set of parameters, the single-particle levels form valley doublets, which become fourfold degenerate when spin is included, and display shell structures reminiscent of a two-dimensional harmonic oscillator with additional lattice-induced splittings. Including Coulomb interactions for two electrons, we find that the ground state remains a triplet over the entire interaction range considered, with a crossover in valley polarization from polarized to unpolarized as the interaction weakens. The corresponding low-energy spectrum, along with its classification by total spin and valley polarization, is presented in the Supplementary Material, which also includes full computational details and analytical considerations.
Motivated by the observation that the $N=2$ ground state remains a triplet across a broad range of interaction strengths, we next examine the case of a double quantum dot (dQD) populated by four electrons. We aim to compare the resulting low-energy spectrum with that of an effective spin-1 Heisenberg model. To this end, we begin by analyzing the single-particle states of the double quantum dot, and then determine the interacting four-electron ground state along with its low-lying excitations.
Figure 2 shows the single-particle spectrum of the double quantum dot for an interdot separation $d=40$ nm. Compared to the single QD case, each shell is now duplicated, with an energy splitting which reflects the presence of another coupled dot. This behaviour can be understood by starting from the single-dot spectrum: as the dots are brought closer together, hybridization between the individual QD states increases, leading to the formation of bonding and antibonding combinations. Consequently, the splitting between these paired states grows as the interdot distance decreases and the depth of the potential increases [45, 46].
Since higher QD states have a more extended wave function, the state hybridization is stronger for higher energy shells. We can see it on the probability density plot for each state (Insets in Fig. 2). Low-energy states like the one in the s-shell are not strongly modified compared to the single QD version. On the other hand, states composed of QD p and d shells are more modified, with the effect being bigger as the state energy increases.
a)Refer to caption b)Refer to caption
Figure 2: (a) Low-energy levels of a double quantum dot. Insets show the corresponding electron densities for each shell, illustrating the increased coupling between interdot states with higher energy, as a consequence of the greater spatial extension of the wavefunctions. (b) Low-energy spectrum for four electrons in a double quantum dot as a function of interaction strength. For weak interactions, the ground state is a valley-unpolarized singlet, while for strong interactions, it becomes a valley-polarized quintuplet. In the regime where the ground state is a singlet, the first and second excited states are a triplet and a quintuplet, respectively, closely resembling the spectrum of a spin-1 Heisenberg antiferromagnetic chain with $L=2$ sites.
Figure 2b shows the low-energy spectrum for four electrons confined in a double quantum dot as a function of interaction strength. In the weak-interaction regime, the ground state is a valley-unpolarized singlet. As the interaction strength increases, the system first transitions to a valley-unpolarized quintuplet ground state and, at even stronger interactions, to a doubly valley-polarized quintuplet. In the singlet ground-state regime, the first and second excited states are a triplet and a quintuplet, respectively. This sequence of spin multiplets closely matches the spectrum of a spin-1 Heisenberg antiferromagnetic chain with $L=2$ sites, indicating that the system’s effective low-energy degrees of freedom behave as emergent spin-1 objects. This correspondence validates the mapping to an effective spin model and underscores the potential of BLGQD arrays for realizing synthetic quantum magnetism. Remarkably, when the ground state becomes a valley-unpolarized quintuplet, the first and second excited states are a triplet and a singlet, respectively, mirroring the spectrum expected for two spin-1s coupled ferromagnetically.
Figure 3: (top) Triplet and quintuplet energies obtained from the fitted BLBQ and Heisenberg models as a function of interaction strength. The blue dashed lines show the energies of the double quantum dot. As discussed in the main text, the BLBQ Hamiltonian accurately reproduces the low-energy spectrum of two coupled quantum dots, each containing two electrons. (bottom) Fitted values of the BLBQ parameters $\beta$ and $J$ as functions of the electron-electron interaction strength.
The bilinear–biquadratic (BLBQ) spin-1 Hamiltonian
$H=J\sum_{i}^{L-1}\left(\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}+\beta\left(\mathbf{S}_{i}\cdot\mathbf{S}_{i+1}\right)^{2}\right)$ (2)
where $\mathbf{S}_{i}$ is the spin operator acting on site $i$, and $J$ is the exchange coupling, $\beta$ is a tunable parameter that controls the strength of the biquadratic interaction, and $L$ is the number of sites. For even $L$, and $\beta<1/3$, the low-energy spectrum consists of a singlet ground state, followed by a triplet and a quintuplet. In particular, for two spins, the energy gaps are given by (see SM for details), $\Delta_{TQ}=2J$, $\Delta_{ST}=J(1-3\beta)$.
Figure 4: (top) Low energies of the $L=100$ BLBQ spin chain as a function of the $\beta$. (bottom) Spin-spin correlation function $g(n)=(-1)^{n}\langle S_{1}^{z}S_{n}^{z}\rangle$ plotted for the first half of the chain. The inset shows the correlation length $\xi$ extracted from exponential fits to $g(n)$, illustrating its dependence on $\beta$. The correlation length $\xi$ characterizes the spatial decay of spin correlations in the system.
Using these expressions, we can map the dQD spectrum onto an effective BLBQ model and extract the corresponding $J$ and $\beta$ parameters. The upper panel of Fig. 3 shows the low energy levels from the fitted BLBQ and Heisenberg models as a function of interaction strength. The blue dashed lines show the energies of the double quantum dot. The BLBQ Hamiltonian reproduces exactly the two low-energy gaps of two coupled quantum dots, containing two electrons each. The lower panel of Fig. 3 shows the fitted value of $J$ and $\beta$ for different values of the ratio between kinetic and Coulomb interacting terms. For stronger interaction, the value of $\beta$ is near zero, as it starts to decrease as the interaction gets weaker. The minimum value shown here for weak interaction is $\beta\approx-0.5$, which is still inside the Haldane phase region.
To explore the behavior of the BLBQ Hamiltonian, we computed the low-energy spectrum of a spin-1 chain with $L=100$ using matrix product state–based density matrix renormalization group (MPS-DMRG) approach [47, 48, 49]. The top panel of Fig. 4 shows the low-energy spectrum as a function of the biquadratic interaction parameter $\beta$ in the range $1/2\leq\beta\leq 0$. In the entire Haldane phase ($1<\beta<1$) and in the thermodynamic limit, the ground state is fourfold degenerate, consisting of a singlet and a triplet manifold separated from the lowest quintuplet by a finite topological gap. This degeneracy arises from two fractionalized spin-$\tfrac{1}{2}$ quasiparticles localized at the chain ends. In a finite chain, the singlet–triplet gap is nonzero but decreases exponentially with system size, reflecting the exponentially small coupling between the edge spins. In contrast, the topological gap to the quintuplet remains finite in the thermodynamic limit and reaches its maximum for $\beta>0$.
The lower panel of Fig. 4 displays the spin-spin correlation function $g(n)=(-1)^{n}\left\langle S_{1}^{z}S_{n}^{z}\right\rangle$ for the first half of the chain. As $\beta$ becomes more negative, the decay of $g(n)$ slows, indicating an increase in the correlation length $\xi$. This length scale is extracted by fitting the envelope of the data to an exponential form, $g(n)\propto\exp(-n/\xi)$. The inset shows the fitted values of $\xi$ as a function of $\beta$. The correlation length $\xi$ controls both the spatial extent of the edge quasiparticles and the range over which they can interact. When $L\gg\xi$, the system is effectively in the thermodynamic limit and the singlet–triplet splitting becomes negligible. However, for $\beta\approx-0.5$, $\xi$ becomes comparable to $L=100$, making finite-size effects visible in the spectrum and enhancing the singlet–triplet splitting. As $\beta\to-1$, the system approaches a quantum critical point characterized by a vanishing energy gap and a diverging correlation length. In this regime, DMRG convergence becomes increasingly challenging due to enhanced entanglement and critical fluctuations.
In summary, we have demonstrated that chains of electrostatically defined bilayer graphene quantum dots (BLGQDs) can become synthetic spin-1 chains, providing an atomistic platform for topological quantum matter. Using a realistic tight-binding description combined with exact diagonalization, we captured Coulomb interactions and valley mixing on equal footing for both single- and double-dot configurations. For two coupled QDs, the low-energy spectrum maps naturally onto a bilinear–biquadratic spin-1 model, establishing a direct link between microscopic electronic states in BLGQDs and effective quantum spin Hamiltonians. This correspondence opens a route to quantum simulators of Haldane topological quantum matter, one-dimensional quantum magnetism with electrically tunable parameters, offering new opportunities for quantum simulation, spintronics, and solid-state quantum information processing.
Acknowledgements.
The author gratefully acknowledges Christoph Stampfer for valuable discussions. This work was supported by the Quantum Sensors Challenge Program QSP-078, High Throughput Networks HTSN-341, and Applied Quantum Computing AQC-004 Challenge Programs at the National Research Council of Canada, NSERC Discovery Grant No. RGPIN-2019-05714, NSERC PQS2D Alliance Quantum Grant No. ALLRP/578466-2022, and University of Ottawa Research Chair in Quantum Theory of Materials, Nanostructures, and Devices. This work was partly enabled by support provided by the Digital Research Alliance of Canada [50].